7:30 – 10:00 VISIBILITY
- When checking over the query logs this morning, found that Paul O. Had spent most of Friday using VISIBILITY. He made 133 queries in 9 or so hours, spending most of his time on the LOTR and cars examples
- Went over status of PPM with Dong. We’re going to aim for Wednesday to walk through the new changes.
10:00 – 3:30 FP
- Working on reading the sensors. They seem to be working well. Working on calibration code – done
- Making a bi-linear normalizing, since there seems to be a big difference between pushing up on a sensor and pushing down.
- Update graphics to show new axis